From GlueXWiki
Jump to: navigation, search

June 5, 2014 FDC+CDC meeting


  1. Gas system [1], [2], [3], [4], [5], Mike's gas GUI
  2. DAQ (Beni)
  3. CDC status (Mike)
  4. Engineering (Bill)
    • cathode measurements [6]
  5. Electronics (Chris, Nick)
  6. Instructions for shift workers (Beni, Lubomir)
  7. FDC alignment [7] (Lubomir)
  8. FDC pedestal resolution: media:pedestal_resolution.pdf (Beni)
  9. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Curtis (CMU), Bill, Beni, Chris, Nick, Simon, Fernando, and Lubomir (JLab).

Gas system

- There were several problems in the FDC parts of the gas system. A series of graphs (from Mike) linked above demonstrate some of them. A 1 l/min leak was found due to bad fitting of one of the valves and later fixed: first graph shows the pressure in the mixing tank (red is FDC) with ~2 l/min filling flow and 0.4l/m FDC usage. To identify the place we were excluding different parts of the system. Once the tank was insulated (flat part in second graph) there was still some pressure drop; it turned out there's a small leak from the relief valve on top of the FDC mixing tank. Same valve is used on the CDC tank but it has not been checked. Narciso is looking for better valves.

- On Tuesday ~6pm there was again a failure of the FDC gas control box in the gas room. As a result the pressure in the mixing tank started dropping below 7 psi. We had to call the guard to open the room and reset the box. A new spare box has been ordered. For the future Beni suggests to have a key from the gas room in the counting house.

- Pressures in the six cells of package 1 are shown in third and fourth graph, when the flow was turned off and then on. It shows very small (~1Pa) variations probably due to temperature and very similar values in the different cells.

- Mike has his first version of the gas GUI ready (screenshot linked above) and checked in the svn: https://halldsvn.jlab.org/repos/trunk/controls/css. Any remarks and recommendations are welcome and should be directed to Mike.

DAQ and CDC status

- No news here. DAQ is in the Sasha's hands now.


- We have a plan to use Fernando's glass scale to do measurements on the cathode. It has to be read out with a microscope with a reticle: see picture done with a cell phone on the top of the eyepiece of the microscope. We can mount a web camera on the eyepiece that has a reticle, or use the built-in microscope camera if it has a software that allows measurements (Bill will check with Chris and Nick about that).


- Fernando discussed the HV issue with CAEN. They say they have higher failure rate for 1550P modules of 3-5 modules per year (we already sent them 6 of these???). In the moment 4 modules are being repaired by CAEN and the turnaround time might be long (2 were sent already beginning of May). Fernando recommended to order 2 spare 1550P modules (each ~$7k). We don't need more spares for 1550N used in the FDC only; we have one spare and 4 in use.

Instructions for shift workers

- For the Hall D readiness review Beni and Lubomir prepared instructions for the shift takers for the CDC,FDC and the Gas System. Look at the link above and give recommendations/corrections to Beni and Lubomir.

FDC alignment

- Lubomir updated the alignment document, still far from final.

FDC pedestal resolution

- Beni demonstrated with the attached plot that the pedestal resolution is improved when you average over more samples. Certainly 6 samples (as it is now) are not enough and we better have 20 or more if the firmware allows it.